OpenStudy (anonymous):

f(x)=[-3x,x<0] [4,x=0] [x^2,x>0] Which statements are true? f(4)=0 f(2)=4 f(3)=9 f(1)=-3 I thought they were all true?

OpenStudy (zzr0ck3r):

for the first one if x = 4 we are looking at thex^2 and 4^2 = 16 for the 3rd one if x = 3 we are lookingat x^2 so 3^2 = 9 for the last one we are lookin at x^2 so 1^2 = 1 so the first, third, and fourth answers are wrong

OpenStudy (zzr0ck3r):

what do we do when x = 2?

OpenStudy (anonymous):

the third answer was right

OpenStudy (zzr0ck3r):

o woops your right

OpenStudy (zzr0ck3r):

do you see why 1 and 4 are wrong?

OpenStudy (anonymous):

hm, not really :C

OpenStudy (anonymous):

Oh, I'm thinking of it wrong ugh but now I got it, was carrying the answer before the comma to the x after the comma and not the = if that makes sense.